OpenStudy (anonymous):

(16^5x)(4^x)= 64 what is x=? could someone please show me how to work this out? thanks

OpenStudy (cggurumanjunath):

16^(5x) * 4^x

OpenStudy (cggurumanjunath):

4^10x*4^x

OpenStudy (cggurumanjunath):

4^11x

OpenStudy (cggurumanjunath):

is the final answer !

OpenStudy (magbak):

do you understand

OpenStudy (cggurumanjunath):

16^5x = 4^2(5x)

OpenStudy (cggurumanjunath):

16=4*4 =4^2

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

you are ment to find x which is the power

OpenStudy (anonymous):